Reduce the quadratic form x1^2 + 2X2x3 to canonical form
1
Expert's answer
2020-10-22T17:58:48-0400

Given quadratic form is "x_1^2+2x_2x_3"


compare the given quadratic expression with standard quadratic form,


"ax_1^2+by_1^2+cz_1^2+2fyz+2gzx+2hxy"


we get,


a=1,b=0,c=0,f=1,g=0 and h=0


Transforming the quadratic equation into matrix form


So the required matrix is ,


"A=\begin{bmatrix}
 a & h &g\\
 h& b&f\\
 g&f&c
 
\end{bmatrix}" , "A=\begin{bmatrix}
 1 & 0 &0\\
 0& 0&1\\
 0&1&0
 
\end{bmatrix}"


Using characterstics Equation "(A-\lambda I)X=0"


To calculate the values of "\lambda"

"|A-\lambda I|=0"


"\begin{vmatrix}
 1 & 0 &0\\
 0& 0&1\\
 0&1&0
 
\end{vmatrix}-\lambda \begin{vmatrix}
 1& 0 &0\\
 0& 1&0\\
 0&0&1
\end{vmatrix}=0"


"\begin{vmatrix}
 1-\lambda & 0 &0\\
 0& -\lambda&1\\
 0&1&-\lambda
 
\end{vmatrix}" =0


Using Factorization method


"(1-\lambda)(\lambda^2-1)=0"


so "\lambda=1,-1,1" (these are the required eigen values)


S0 the required canonical form is


"\lambda_1x^2+\lambda_2y^2+\lambda_3z^2=0"


"x^2-y^2+z^2=0"

This is the required canonical form.